Equipment acquired at the beginning of the fiscal year at a cost of $144,000 has an estimated residual value of $9,000 and an estimated useful life of 10 years.

a. Determine the amount of annual depreciation by the straight-line method. $

b. Determine the amount of depreciation for the first and second years computed by the double-declining-balance method.
Answer #1

a.Annual depreciation=(Cost-Residual Value)/Useful Life

=(144,000-9000)/10=$13500

b.Annual depreciation rate as per straight line method=100%/10=10%/year

Hence depreciation rate as per double decline=2*Annual depreciation rate as per straight line method*Beginning value of each period

Year Beginning value Depreciation Ending value
1 144,000 (2*10%*144,000)=$28800. (144,000-28800)=$115200
2 115200 (2*10%*115200)=$23040.